4. Find the square roots of the following numbers by the Prime Factorisation Method.  iii. 1764  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


1764 = 2×2×3×3×7×7
⇒ 1764 = (2×2)×(3×3)×(7×7) ⇒ 1764 = (2×3×7)×(2×3×7) ⇒ 1764 = (2×3×7)2⇒√1764=2×3×7=42